OpenStudy (anonymous):

for the function f(x) = ln x/x^2, find the approximate location of the critical point in the interval (0,5). (How to get the second x value) so far I have (1.6487,x)

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve for \(x\) you got the derivative?

OpenStudy (anonymous):

(1-2 log(x))/ (x^20

OpenStudy (anonymous):

i assume you mean \[\frac{1-2\ln(x)}{x^3}\]

OpenStudy (anonymous):

set \[1-2\ln(x)=0\] and solve for \(x\) to find your critical point

OpenStudy (anonymous):

you good with that?

OpenStudy (anonymous):

Hmm yes. I haven't solved yet, but for x I should have two values correct?

OpenStudy (anonymous):

I got 1.648, but there are suppose to be two x's. How?

OpenStudy (anonymous):

i get \[\sqrt{e}\]

OpenStudy (anonymous):

there are not two, not sure why there should be

OpenStudy (anonymous):

Ohh I've found why, I must plug in the 1.648 into the original equation, so it reads ln(1.649)/1.649^2.